The weights of 1000 children, in average is 61kg with a standard deviation of 18kg. Suppose the weights ate normally distributed, how many children weigh less than 35kg?


1
Expert's answer
2021-06-03T11:40:17-0400

Solution:

Given, "\mu=61,\sigma=18"

"X\sim N(\mu,\sigma)
\\P(X<35)=P(z<\dfrac{35-61}{18})=P(z<-1.44)=0.07493"

Number of required children"=1000\times0.07493=74.93\approx75"